# ISDrama: Immersive Spatial Drama Generation through Multimodal Prompting

### v1

@@ Abstract
Multimodal immersive spatial drama generation focuses on creating continuous multi-speaker binaural speech with dramatic prosody based on multimodal prompts, with potential applications in AR, VR, and others. This task requires simultaneous modeling of spatial information and dramatic prosody based on multimodal inputs, with high data collection costs. To the best of our knowledge, our work is the first attempt to address these challenges. We construct MRSDrama, the first multimodal recorded spatial drama dataset, containing binaural drama audios, scripts, videos, geometric poses, and textual prompts. Then, we propose ISDrama, the first immersive spatial drama generation model through multimodal prompting. ISDrama comprises these primary components: 1) Multimodal Pose Encoder, based on contrastive learning, considering the Doppler effect caused by moving speakers to extract unified pose information from multimodal prompts. 2) Immersive Drama Transformer, a flow-based mamba-transformer model that generates high-quality drama, incorporating Drama-MOE to select proper experts for enhanced prosody and pose control. We also design a context-consistent classifier-free guidance strategy to coherently generate complete drama. Experimental results show that ISDrama outperforms baseline models on objective and subjective metrics. The demos and dataset are available at https://aaronz345.github.io/ISDramaDemo.
Machine Learning, ICML
1 Introduction
@@
Figure 1: ISDrama uses scripts as content, prompt audio to guide timbre, and spatial information from multimodal prompts, to create continuous multi-speaker binaural speech with dramatic prosody.
@@
Binaural hearing provides localization cues through sound fields, enhancing the human spatial perception of the environment. This capability is critical for applications requiring deep immersion, like movies, VR, and AR (Fitria, 2023). Compared to non-linguistic binaural audio (Sun et al., 2024), generating binaural speech is more challenging yet promising. Specifically, based on multimodal prompts from diverse contexts, generating continuous multi-speaker binaural speech with dramatic prosody can create immersive spatial drama. This new task enhances storytelling, offering an immersive emotional and spatial engagement, and an integration of virtual and reality.
@@
With deep learning advancements, progress has been made in synthesizing speech with prosody modeling (Ju et al., 2024; Du et al., 2024) and generating binaural audio from monaural audio based on input multimodal prompts (Li et al., 2024d; Parida et al., 2022). Cascading these methods to generate continuous multi-speaker binaural speech with prosody modeling seems to be a viable approach. However, this cascading method will disrupt the integration of prosody and spatial modeling, leading to mismatches and error accumulation. This highlights the need for a unified multimodal immersive spatial drama generation framework.
@@
To address these challenges, we first introduce MRSDrama, the first multimodal recorded spatial drama dataset, comprising binaural drama audios, scripts, videos, geometric poses, and textual prompts. The dataset includes 97.82 hours of speech data recorded by 21 speakers across three scenes. Next, we propose ISDrama, the first immersive spatial drama generation model based on multimodal prompting. ISDrama generates high-quality, continuous, multi-speaker binaural speech with dramatic prosody and spatial immersion, driven by multimodal prompts. To extract a unified pose representation from multimodal prompts, we design the Multimodal Pose Encoder, a contrastive learning-based framework that encodes not only position and head orientation but also radial velocity, accounting for the Doppler effect caused by moving speakers. Meanwhile, we develop the Immersive Drama Transformer, a flow-based Mamba-Transformer model capable of generating immersive spatial drama effectively and stably. Within this model, we introduce Drama-MOE (Mixture of Experts), which selects the appropriate experts to enhance prosodic expressiveness and improve pose control. Then, we adopt a context-consistent classifier-free guidance (CFG) strategy to ensure the quality and coherence of complete drama generation. We evaluate ISDrama on quality, speaker similarity, prosodic expressiveness, pose, angle, distance, IPD, and ILD. Experimental results show that ISDrama outperforms baseline models on both objective and subjective metrics, demonstrating its ability to generate immersive spatial drama that adheres to physical principles while exhibiting rich prosodic variation. Overall, our main contributions can be summarized as:
@@
  • We introduce ISDrama, the first immersive spatial drama generation model through multimodal prompting. We design the Multimodal Pose Encoder to extract pose from multimodal inputs, while the Immersive Drama Transformer produces binaural speech.
@@
Table 1: Comparison of current open-source recorded spatial speech datasets. Multi-channel speech does not account for the intricate structures of human ears, whereas binaural speech incorporates natural IPD and ILD, ensuring a realistic and immersive hearing perception.
@@
Figure 2: The pipeline of MRSDrama data collection. Human double-checks exist in each process. Notably, all data are desensitized.
2 Related Work
Audio Spatialization with Multimodal Cues. In recent years, the development of deep learning has significantly advanced the exploration of spatial audio. Substantial progress has been made in sound source localization for binaural audio (Krause et al., 2023; Yang & Zheng, 2022; Shimada et al., 2022; García-Barrios et al., 2022). At the same time, the rise of multimodal research has spurred innovations in spatial audio synthesis. Mono2Binaural (Gao & Grauman, 2019) devises a deep convolutional neural network that leverages visual cues to convert monaural audio into binaural audio. Sep-Stereo (Zhou et al., 2020) enhances stereo audio by integrating audio-visual features. Garg et al. (2021) proposes a multi-task framework that learns geometry-aware features for generating binaural audio. CLUP (Li et al., 2024d) jointly learns to localize visually sounding objects and generate binaural audio. Beyond visual modalities, BinauralGrad (Leng et al., 2022) employs positional information for monaural-to-stereo audio generation, while scene depth maps have also been utilized (Parida et al., 2022). However, these methods rely on monaural audio input, limiting their applicability for complex and flexible generation tasks. Recently, SpatialSonic (Sun et al., 2024) employs a Diffusion Transformer to generate binaural audio from both text and image prompts. VISAGE (Kim et al., 2025) leverages CLIP visual features to generate first-order ambisonics directly from silent video frames. However, these tasks focus on generating short audio clips lacking actual linguistic information. ImmerseDiffusion (Heydari et al., 2025) proposes an end-to-end generative audio model conditioned on textual prompts of spatial, temporal, and environmental factors. However, it lacks support for multimodal prompts, binaural scenes, and expressive prosody modeling. Generating continuous linguistic speech with unified scene and pose information extracted from more diverse prompt modalities remains an unresolved challenge.
@@ 2 Related Work
Speech Synthesis with Prosody Modeling. Prosody modeling aims to synthesize target speech with natural or emotional prosody, which is essential to generate expressive speech in a controlled manner and typically involves transferring prosody from prompt audio (Wagner & Watson, 2010). Skerry-Ryan et al. (2018) is the first to integrate a prosody reference encoder into a Tacotron-based TTS system, enabling the transfer of prosody for similar-text speech. Attentron (Choi et al., 2020) introduces an attention mechanism to extract prosody from reference samples. ZSM-SS (Kumar et al., 2021) proposes a Transformer-based architecture with an external speaker encoder using wav2vec 2.0 (Baevski et al., 2020). Li et al. (2021) incorporates global utterance-level and local phoneme-level prosodic features in target speech. Daft-Exprt (Zaıdi et al., 2021) employs a gradient reversal layer to enhance target speaker fidelity in prosody transfer. Generspeech (Huang et al., 2022) incorporates the attention mechanism to capture multi-level prosody. HierSpeech++ (Lee et al., 2023) generates F0 representation based on text representations and prosody prompts, while StyleTTS 2 (Li et al., 2024c) predicts pitch and energy based on a prosody predictor (Li et al., 2022). Mega-TTS 2 (Jiang et al., 2024) employs clustering VQ for prosody encoding and Prosody Latent LM for predicting. NaturalSpeech 3 (Ju et al., 2024) employs factorized vector quantization to disentangle prosody. CosyVoice (Du et al., 2024) incorporates x-vectors into an LLM to disentangle and model prosody. FireRedTTS (Guo et al., 2024) employs a semantic-aware speech tokenizer to encode speech style and conduct style control. However, these models only focus on monaural speech synthesis and cannot simultaneously model dramatic prosody, which requires semantic alignment in the temporal-frequency domain, and spatial information, which spans the spatial-temporal dimensions.
Figure 3: The statistics of MRSDrama. The position distribution is plotted on the plane defined by the listener’s forward direction and ears.
3 Dataset: MRSDrama
@@
Annotation. We separately annotate the audio and video components. 1) For Audio Annotation, We first perform denoising using FRCRN (Zhao et al., 2022). Subsequently, MFA (McAuliffe et al., 2017) is employed for a coarse phoneme-to-audio alignment between the original script and audio. Chinese phonemes are extracted by pypinyin. Next, annotators are asked to use Praat (Boersma, 2001) to refine the rough alignment, focusing on correcting word and phoneme boundaries and addressing erroneous words based on auditory perception. 2) For Video Annotation, Speakers occasionally move between many fixed points in each scene. Annotators are asked to record speakers’ arrival times and point coordinates of each movement. They then measure these speakers’ head orientation and mouth height while standing and sitting in each frame, extracting 3D position coordinates and quaternion orientation to form frame-level sound source poses. Based on the annotated pose data, we then use GPT-4o (Achiam et al., 2023) to generate textual prompts for each actor’s line by combining the orientation, endpoint, direction, speed, and start time of each motion. Meanwhile, annotators also label the camera pose (position and orientation) and scene prompts, including descriptions about room sizes and acoustical effects.
@@
4 Method: ISDrama
Figure 4: The architecture of ISDrama. In Figure (a), DP & LR means duration predictor and length regulator. In Figure (b), the ODE Solver, autoencoder decoder, and vocoder generate predicted audio from Gaussian noise with conditions during inference. In Figure (c), MPE is the Multimodal Pose Encoder, while IDT denotes the Immersive Drama Transformer. MPE predicts content duration and segments the inputs. Then IDL coherently generates the complete drama. In Figure (d), FAN denotes Fourier Analysis Networks.
4.1 Overview
@@
4.2 Multimodal Pose Encoder
@@
The correct phase estimation is crucial for binaural audio (Richard et al., 2022). Therefore, for geometric pose, we not only encode the head orientation $ori$ and 3D coordinates of the sound source relative to left and right ears, $\vec{pos}_{l}$ and $\vec{pos}_{r}$ but also add the radial relative velocity according to Doppler effect (Gill, 1965) for phase estimation (Liu et al., 2022a). Specifically, we calculate the 3D velocity vector $\vec{v}$ of the moving sound source in the Cartesian coordinate system, then decompose $\vec{v}$ into radial velocity components $v_{rad-l}$ and $v_{rad-r}$ in the spherical coordinate systems of left and right ears respectively by $\vec{v}_{rad}=\frac{\vec{pos}\cdot\vec{v}}{\|\vec{pos}\|}\cdot\hat{\mathbf{r}}$, where $\hat{\mathbf{r}}\in\mathcal{R}^{1}$ is the radial unit vector. Finally, we can encode and concatenate $(\vec{pos}_{l},\vec{pos}_{r},ori,v_{rad-l},v_{rad-r})$ into $z_{p-geo}$.
To determine the target embedding length and considering the relationship between semantics, spoken speed, and motion, we encode the pronunciation and semantics of content $c$ and expand it with the predicted phoneme duration as $z_{c}$. Since the input script is organized in speaker transitions, we can compute predicted phoneme durations to segment the length for each target speech as shown in Figure 4 (c). Then we employ a Transformer to predict $z_{p-txt}$ and $z_{p-vid}$, where $z_{c}$ serves as the input and leverages different modal conditions. For textual prompts, we use FLAN-T5 (Chung et al., 2024) to encode text as the condition. For silent video, we concatenate camera pose, mouth pixel sequences from Co-Tracker3 (Karaev et al., 2024), and embedding from CLIP (Radford et al., 2021) as the condition.
@@
$\mathcal{L}_{p1,p2}=-\frac{1}{2N}\sum_{i=1}^{N}(\log\frac{\exp({sim}({z_{p1}}^{i},{z_{p2}}^{i})/\tau)}{\sum_{j=1}^{N}\exp({sim}({z_{p1}}^{i},{z_{p2}}^{j})/\tau)}$ (1)
@@
4.3 Immersive Drama Transfomer
Flow-based Mamba-Transformer. Spatial information and prosody in binaural speech are related to pose, scene, emotion, and rhythm. Their interaction necessitates the simultaneous modeling of both temporal-frequency and spatial-temporal, which becomes more complex with long sequences. The flow matching method facilitates smooth transformations, promoting stable and rapid generation, while the Transformer architecture excels in sequence generation, making the flow-based Transformer particularly suitable for this task. Meanwhile, since Transformer tends to be computationally expensive for long sequences, we add Mamba blocks (Gu & Dao, 2023) at an Attention-to-Mamba ratio of 1:K to balance memory usage, efficient training, and the ability to handle long sequences.
As shown in Figure 4 (b), we add Gaussian noise $\epsilon$ to the autoencoder output $\hat{m_{gt}}$ to obtain $x_{t}$ at timestep $t$. We then add content embedding $z_{c}$ and pose embedding $z_{p}$ to $x_{t}$ and concatenate it with prompt audio embedding $z_{a}$. Therefore, using the self-attention mechanism and capability of Mamba blocks to capture long-range dependencies, Immersive Drama Transformer can effectively model content, pose, timbre, accent, and articulation. Scene information $s$ (e.g., a video frame in silent video or a textual description for other inputs) is encoded and processed using the cross-attention mechanism to simulate the room’s acoustic properties, such as the differences in acoustic effects caused by varying room sizes and acoustical effects. Notably, we use the first block’s output to predict binaural F0, providing extra supervision and additional input for subsequent blocks, helping to model dramatic prosody. Furthermore, we employ RMSNorm (Zhang & Sennrich, 2019) and design global adapters with AdaLN (Peebles & Xie, 2023) to ensure training stability and global consistency in timbre and scene. The final output vector field is trained with the flow-matching objective:
$\mathcal{L}_{flow}=\mathbb{E}_{t,p_{t}(x_{t})}\left\|v_{t}(x_{t},t|C;\theta)-(\hat{m_{gt}}-\epsilon)\right\|^{2},$ (2)

### v6

@@ Abstract.
Multimodal immersive spatial drama generation focuses on creating continuous multi-speaker binaural speech with dramatic prosody based on multimodal prompts, with potential applications in AR, VR, and others. This task requires simultaneous modeling of spatial information and dramatic prosody based on multimodal inputs, with high data collection costs. To the best of our knowledge, our work is the first attempt to address these challenges. We construct MRSDrama, the first multimodal recorded spatial drama dataset, containing binaural drama audios, scripts, videos, geometric poses, and textual prompts. Then, we propose ISDrama, the first immersive spatial drama generation model through multimodal prompting. ISDrama comprises these primary components: 1) Multimodal Pose Encoder, based on contrastive learning, considering the Doppler effect caused by moving speakers to extract unified pose information from multimodal prompts. 2) Immersive Drama Transformer, a flow-based mamba-transformer model that generates high-quality drama, incorporating Drama-MOE to select proper experts for enhanced prosody and pose control. We also design a context-consistent classifier-free guidance strategy to coherently generate complete drama. Experimental results show that ISDrama outperforms baseline models on objective and subjective metrics. The demos are available at https://aaronz345.github.io/ISDramaDemo. We provide the dataset and the evaluation code at https://huggingface.co/datasets/AaronZ345/MRSDrama and https://github.com/AaronZ345/ISDrama.
immersive spatial drama generation, spatial audio dataset, prosody modeling, speech synthesis, multimodal prompting
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755014isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Artificial intelligenceccs: Applied computing Arts and humanities
1. Introduction
@@
Figure 1. ISDrama uses scripts as content, prompt audio to guide timbre, and spatial information from multimodal prompts, to create continuous multi-speaker binaural speech with dramatic prosody.
[…]
ISDrama uses scripts as content, prompt audio to guide timbre, and spatial information from multimodal prompts, to create continuous multi-speaker binaural speech with dramatic prosody.
@@
Binaural hearing provides localization cues through sound fields, enhancing the human spatial perception of the environment. This capability is critical for applications requiring deep immersion, like movies, VR, and AR (Fitria, 2023; Huiyu et al., 2025; Bosun, 2020). Compared to non-linguistic binaural audio (Sun et al., 2024), generating binaural speech is more challenging yet promising. Specifically, based on multimodal prompts from diverse contexts, generating continuous multi-speaker binaural speech with dramatic prosody can create immersive spatial drama. This new task enhances storytelling, offering an immersive emotional and spatial engagement, and an integration of virtual and reality.
@@
With deep learning advancements, progress has been made in synthesizing speech with prosody modeling (Ju et al., 2024; Du et al., 2024) and generating binaural audio from monaural audio based on input multimodal prompts (Li et al., 2024d; Parida et al., 2022). Cascading these methods to generate continuous multi-speaker binaural speech with prosody modeling seems to be a viable approach. However, this cascading method will disrupt the integration of prosody and spatial modeling, leading to mismatches and error accumulation. This highlights the necessity for a unified framework for multimodal immersive spatial drama generation.
@@
To address these challenges, we first introduce MRSDrama, the first multimodal recorded spatial drama dataset, comprising binaural drama audios, scripts, videos, geometric poses, and textual prompts. The dataset includes 97.82 hours of speech data recorded by 21 speakers across three scenes. Next, we propose ISDrama, the first immersive spatial drama generation model based on multimodal prompting. ISDrama generates high-quality, continuous, multi-speaker binaural speech with dramatic prosody and spatial immersion, driven by multimodal prompts. To extract a unified pose representation from multimodal prompts, we design the Multimodal Pose Encoder, a contrastive learning-based framework that encodes not only position and head orientation but also radial velocity, accounting for the Doppler effect caused by moving speakers. Meanwhile, we develop the Immersive Drama Transformer, a flow-based Mamba-Transformer model capable of generating immersive spatial drama effectively and stably. Within this model, we introduce Drama-MOE (Mixture of Experts), which selects the appropriate experts to enhance prosodic expressiveness and improve pose control. Then, we adopt a context-consistent classifier-free guidance (CFG) strategy to ensure the quality and coherence of complete drama generation. We evaluate ISDrama on quality, speaker similarity, prosodic expressiveness, pose, angle, distance, IPD, and ILD. Experimental results show that ISDrama outperforms baseline models on both objective and subjective metrics, demonstrating its ability to generate immersive spatial drama that adheres to physical principles while exhibiting rich prosodic variation. Overall, our main contributions can be summarized as follows:
@@
  • We introduce ISDrama, the first immersive spatial drama generation model through multimodal prompting. We design the Multimodal Pose Encoder to extract pose information from multimodal inputs, while the Immersive Drama Transformer produces high-quality binaural speech.
@@
2. Related Work
Audio Spatialization with Multimodal Cues. In recent years, the development of deep learning has significantly advanced the exploration of spatial audio. Substantial progress has been made in sound source localization for binaural audio (Krause et al., 2023; Yang and Zheng, 2022; Shimada et al., 2022; Garc´ıa-Barrios et al., 2022). At the same time, the rise of multimodal research has spurred innovations in spatial audio synthesis. Mono2Binaural (Gao and Grauman, 2019) devises a deep convolutional neural network that leverages visual cues to convert monaural audio into binaural audio. Sep-Stereo (Zhou et al., 2020) enhances stereo audio by integrating audio-visual features. Garg et al. (2021) proposes a multi-task framework that learns geometry-aware features for generating binaural audio. CLUP (Li et al., 2024d) jointly learns to localize visually sounding objects and generate binaural audio. Beyond visual modalities, BinauralGrad (Leng et al., 2022) employs positional information for monaural-to-stereo audio generation, while scene depth maps have also been utilized (Parida et al., 2022). However, these methods rely on monaural audio input, limiting their applicability for complex and flexible generation tasks. Recently, SpatialSonic (Sun et al., 2024) employs a Diffusion Transformer to generate binaural audio from both text and image prompts. VISAGE (Kim et al., 2025) leverages CLIP visual features to generate first-order ambisonics directly from silent video frames. However, these tasks focus on generating short audio clips lacking actual linguistic information. ImmerseDiffusion (Heydari et al., 2025) proposes an end-to-end generative audio model conditioned on textual prompts of spatial, temporal, and environmental factors. However, it lacks support for multimodal prompts, binaural scenes, and expressive prosody modeling. Generating continuous linguistic speech with unified scene and pose information extracted from more diverse prompt modalities remains an unresolved challenge.
Table 1. Comparison of open-source recorded spatial speech datasets. Multi-channel speech does not account for the intricate structures of human ears, whereas binaural speech incorporates natural IPD and ILD, ensuring an immersive hearing perception.
@@ 2. Related Work
Figure 2. The pipeline of MRSDrama data collection. Human double-checks exist in each process. All data are desensitized.
[…]
The pipeline of MRSDrama data collection.
@@
Speech Synthesis with Prosody Modeling. Prosody modeling aims to synthesize target speech with natural or emotional prosody, which is essential to generate expressive speech in a controlled manner and typically involves transferring prosody from prompt audio (Wagner and Watson, 2010). Skerry-Ryan et al. (2018) is the first to integrate a prosody reference encoder into a Tacotron-based TTS system, enabling the transfer of prosody for similar-text speech. Attentron (Choi et al., 2020) introduces an attention mechanism to extract prosody from reference samples. ZSM-SS (Kumar et al., 2021) proposes a Transformer-based architecture with an external speaker encoder using wav2vec 2.0 (Baevski et al., 2020). Li et al. (2021) incorporates global utterance-level and local phoneme-level prosodic features in target speech. Daft-Exprt (Zaıdi et al., 2021) employs a gradient reversal layer to enhance target speaker fidelity in prosody transfer. Generspeech (Huang et al., 2022) incorporates the attention mechanism to capture multi-level prosody. HierSpeech++ (Lee et al., 2023) generates F0 representation based on text representations and prosody prompts, while StyleTTS 2 (Li et al., 2024a) predicts pitch and energy based on a prosody predictor (Li et al., 2022). NaturalSpeech 3 (Ju et al., 2024) employs factorized vector quantization to disentangle prosody. CosyVoice (Du et al., 2024) incorporates x-vectors into an LLM to disentangle and model prosody. FireRedTTS (Guo et al., 2024) employs a semantic-aware speech tokenizer to encode speech style. However, these models only focus on monaural speech synthesis and cannot simultaneously model dramatic prosody, which requires semantic alignment in the temporal-frequency domain, and spatial information, which spans the spatial-temporal dimensions.
3. Dataset: MRSDrama
@@
Figure 3. The statistics of MRSDrama. The position distribution is plotted on the plane defined by listeners’ direction and ears.
[…]
The statistics of MRSDrama.
Annotation. We separately annotate the audio and video components. 1) For Audio Annotation, We first perform denoising using FRCRN (Zhao et al., 2022). Subsequently, MFA (McAuliffe et al., 2017) is employed for a coarse phoneme-to-audio alignment between the original script and audio. Chinese phonemes are extracted by pypinyin. Next, annotators are asked to use Praat (Boersma, 2001) to refine the rough alignment, focusing on correcting word and phoneme boundaries and addressing erroneous words based on auditory perception. 2) For Video Annotation, Speakers occasionally move between many fixed points in each scene. Annotators are asked to record speakers’ arrival times and point coordinates of each movement. They then measure these speakers’ head orientation and mouth height while standing and sitting in each frame, extracting 3D position coordinates and quaternion orientation to form frame-level sound source poses. Based on the annotated pose data, we then use GPT-4o (Achiam et al., 2023) to generate textual prompts for each actor’s line by combining the orientation, endpoint, direction, speed, and start time of each motion. Meanwhile, annotators also label the camera pose (position and orientation) and scene prompts, including room sizes and acoustical effects.
@@
4. Method: ISDrama
Figure 4. The architecture of ISDrama. In Figure (a), DP & LR is duration predictor and length regulator. In Figure (b), the ODE Solver, autoencoder decoder, and vocoder generate predicted audio from Gaussian noise with conditions during inference. In Figure (c), MPE is the Multimodal Pose Encoder, while IDT is the Immersive Drama Transformer. MPE predicts content duration and segments inputs. Then IDL coherently generates the complete drama. In Figure (d), FAN is Fourier Analysis Networks.
[…]
The architecture of ISDrama.
4.1. Overview
@@
4.2. Multimodal Pose Encoder
@@
The correct phase estimation is crucial for binaural audio (Richard et al., 2022). Therefore, for geometric pose, we not only encode the head orientation $ori$ and 3D coordinates of the sound source relative to left and right ears, $\vec{pos}_{l}$ and $\vec{pos}_{r}$ but also add the radial relative velocity according to Doppler effect (Gill, 1965) for phase estimation (Liu et al., 2022b). Specifically, we calculate the 3D velocity vector $\vec{v}$ of the moving sound source in the Cartesian coordinate system, then decompose $\vec{v}$ into radial velocity components $v_{rad-l}$ and $v_{rad-r}$ in the spherical coordinate systems of left and right ears respectively by $\vec{v}_{rad}=\frac{\vec{pos}\cdot\vec{v}}{\|\vec{pos}\|}\cdot\hat{\mathbf{r}}$, where $\hat{\mathbf{r}}\in\mathcal{R}^{1}$ is the radial unit vector. Finally, we can encode and concatenate $(\vec{pos}_{l},\vec{pos}_{r},ori,v_{rad-l},v_{rad-r})$ into $z_{p-geo}$.
To determine the target embedding length and considering the relationship between semantics, spoken speed, and motion, we encode the pronunciation and semantics of content $c$ and expand it with the predicted phoneme duration as $z_{c}$. Since the input script is organized in speaker transitions, we can compute predicted phoneme durations to segment the length for each target speech as shown in Figure 4 (c). Then we employ a Transformer to predict $z_{p-txt}$ and $z_{p-vid}$, where $z_{c}$ serves as the input and leverages different modal conditions. For textual prompts, we use FLAN-T5 (Chung et al., 2024) to encode text as the condition. For silent video, we encode and concatenate camera pose, mouth pixel sequences from Co-Tracker3 (Karaev et al., 2024), and embedding from CLIP (Radford et al., 2021) as the condition.
@@
(1) $\mathcal{L}_{p1,p2}=-\frac{1}{2N}\sum_{i=1}^{N}(\log\frac{\exp({sim}({z_{p1}}^{i},{z_{p2}}^{i})/\tau)}{\sum_{j=1}^{N}\exp({sim}({z_{p1}}^{i},{z_{p2}}^{j})/\tau)}$
@@
4.3. Immersive Drama Transfomer
Flow-based Mamba-Transformer. Spatial information and prosody in binaural speech are related to pose, scene, emotion, and rhythm. Their interaction necessitates the simultaneous modeling of both temporal-frequency and spatial-temporal, which becomes more complex with long sequences. The flow matching method facilitates smooth transformations, promoting stable and rapid generation, while the Transformer architecture excels in sequence generation, making the flow-based Transformer particularly suitable for this task. Meanwhile, since Transformer tends to be computationally expensive for long sequences, we add Mamba blocks (Gu and Dao, 2023) at an Attention-to-Mamba ratio of 1:K to balance memory usage, efficient training, and the ability to handle long sequences.
As shown in Figure 4 (b), we add Gaussian noise $\epsilon$ to the autoencoder output $\hat{m_{gt}}$ to obtain $x_{t}$ at timestep $t$. We then add content embedding $z_{c}$ and pose embedding $z_{p}$ to $x_{t}$ and concatenate it with prompt audio embedding $z_{a}$. Therefore, using the self-attention mechanism and capability of Mamba blocks to capture long-range dependencies, Immersive Drama Transformer can effectively model content, pose, timbre, accent, and articulation. Scene information $s$ (e.g., a video frame in silent video or a textual description for other inputs) is encoded and processed using the cross-attention mechanism to simulate the room’s acoustic properties, such as the differences in acoustic effects caused by varying room sizes and acoustical effects. Notably, we use the first block’s output to predict binaural F0, providing extra supervision and additional input for subsequent blocks, helping to model dramatic prosody. Furthermore, we employ RMSNorm (Zhang and Sennrich, 2019) and design global adapters with AdaLN (Peebles and Xie, 2023) to ensure training stability and global consistency in timbre and scene. The final output vector field is trained as:
(2) $\mathcal{L}_{flow}=\mathbb{E}_{t,p_{t}(x_{t})}\left\|v_{t}(x_{t},t|C;\theta)-(\hat{m_{gt}}-\epsilon)\right\|^{2},$